OpenStudy (anonymous):

Simplifysquare root of 72 end root minus square root of 147 end root minus 5 square root of 8 end root plus 3 square root of 48 negative 2 square root of 35 square root of 5 negative 8 plus 5 square root of 6 negative 4 square root of 2 end root plus 5 square root of 3

OpenStudy (anonymous):

\[\sqrt{72} - \sqrt{147} - 5\sqrt{8} + 3\sqrt{48}\]

OpenStudy (anonymous):

A) \[-2\sqrt{35}\]

OpenStudy (anonymous):

B) \[\sqrt{5}\]

OpenStudy (anonymous):

C) \[-8 + 5\sqrt{6}\]

OpenStudy (anonymous):

D) \[-4\sqrt{2} + 5\sqrt{3}\]

OpenStudy (anonymous):

@lacrosseplayer22 @dan815 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Hint: \[\Large \sqrt{72} = \sqrt{36*2}\] \[\Large \sqrt{72} = \sqrt{36}*\sqrt{2}\] \[\Large \sqrt{72} = 6*\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

Simplify the other roots the same way

OpenStudy (anonymous):

what do you do after all the roots are simplified? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

You combine like terms

OpenStudy (anonymous):

\[6\sqrt{2} - 7\sqrt{3} - 10\sqrt{2} + 12\sqrt{3}\]

jimthompson5910 (jim_thompson5910):

good, you've simplified correctly now combine like terms

OpenStudy (anonymous):

\[-4\sqrt{2} + 5\sqrt{3}\]

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thank you!!

jimthompson5910 (jim_thompson5910):

you're welcome